OpenStudy (anonymous):

What is a simpler form of each product? (4x – 6y3)^2

OpenStudy (anonymous):

\[\large {(a-b)^2=a^2-2ab+b^2}\]

OpenStudy (anonymous):

(4x – 6y3)^2 = (4x – 6y3) * (4x – 6y3) Now just use the foil method

OpenStudy (anonymous):

It's the same thing you have been doing for the last two problems

OpenStudy (anonymous):

i keep getting a really big number.

OpenStudy (anonymous):

Post it ,We'll check it

OpenStudy (anonymous):

well for 6^3 im getting 216.

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

okay, what do i do next?

OpenStudy (anonymous):

Where is \( 6^3 \)

OpenStudy (anonymous):

16x^2-48xy^3+36y^6

OpenStudy (anonymous):

\[ (4x – 6y^3) * (4x – 6y^3) =4x*4x+ (-6y^3)*4x+4x*(-6y^3)+(-6y^3)*(-6y^3)\]

OpenStudy (anonymous):

This is vat u get ven u expand.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

U r welcum